OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now?

OpenStudy (e.cociuba):

Let Sam's age=x Sam's coach age=2x ____________________________ 10 years before, Sam's age=x-10 10 years before, Sam's coach age=3(x-10)=3x-30 2x-10=3x-30 3x-2x=-10+30 ______________________ x=20 yrs old

OpenStudy (anonymous):

Thank,You

OpenStudy (e.cociuba):

No probelmo! :)